# RobustBlack: Challenging Black-Box Adversarial Attacks on State-of-the-Art Defenses

### v1

@@ Abstract
Although adversarial robustness has been extensively studied in white-box settings, recent advances in black-box attacks (including transfer- and query-based approaches) are primarily benchmarked against weak defenses, leaving a significant gap in the evaluation of their effectiveness against more recent and moderate robust models (e.g., those featured in the Robustbench leaderboard). In this paper, we question this lack of attention from black-box attacks to robust models. We establish a framework to evaluate the effectiveness of recent black-box attacks against both top-performing and standard defense mechanisms, on the ImageNet dataset. Our empirical evaluation reveals the following key findings: (1) the most advanced black-box attacks struggle to succeed even against simple adversarially trained models; (2) robust models that are optimized to withstand strong white-box attacks, such as AutoAttack, also exhibits enhanced resilience against black-box attacks; and (3) robustness alignment between the surrogate models and the target model plays a key factor in the success rate of transfer-based attacks.
1 Introduction
@@
Since the discovery that deep learning models are susceptible to minor input disturbances, yielding adversarial examples (AE) (Goodfellow et al., 2015), the development of robust models has become one of the most active topics in the machine learning community. This topic has been thoroughly explored for white-box settings, leading to standardized attacks (Croce & Hein, 2020b) and benchmarks (Croce et al., 2020).
However, realistic evaluations of adversarial attacks remain less explored. It involves scenarios where the attacker has limited or no knowledge of the target model to be attacked. These scenarios are referred to as black-box (or gray-box with partial knowledge (Guo et al., 2018)). Black-box attacks belong to one of the two following settings: When the output of the model provides both category labels and scores, it is referred to as a score-based black-box attack. Conversely, if the output includes only category labels, it is named a decision-based black-box attack (Brendel et al., 2018). Given that most computer vision Application Programming Interface (APIs) (Google, 2023; Imagga, 2023) generally provide image prediction categories alongside scores, most research in a black-box setting considers score-based black-box attacks. Hence, this work focuses on this category.
@@
State of the art (SOTA) black-box attacks leverage one or both of the following mechanisms: (1) adversarial examples transferability and (2) iterative queries with meta-heuristics. Transferability methods craft adversarial examples on a specifically built surrogate model (with white-box access) such that this example should transfer to (i.e. also fools) a target model (to which they have black-box access). Iterative query methods use query feedback from the target model to guide or refine the optimization process to uncover its vulnerabilities.
Transferability is difficult to achieve when the target differs from the surrogate (in terms of architecture, training, or dataset) and the factors behind transferability (or lack thereof) remain an active field of study (Charles et al., 2020; Dong et al., 2018a; Li et al., 2018; Wu et al., 2020a). This is because adversarial attacks crafted on the surrogate lead to examples that maximize the loss function of the surrogate model (Goodfellow et al., 2015; Kurakin et al., 2017), while the target model has a different loss function. Methods to improve transferability mostly rely on building diversity during the optimization of adversarial examples (Li et al., 2018; Wu et al., 2020a; Xie et al., 2019a).
Iterative-query attacks also face multiple challenges: With increased search space (large images, for example), common query attacks show a significant decrease in the Attack Success Rate (ASR) or require large amounts of queries to achieve sufficient ASR (Mohaghegh Dolatabadi et al., 2020; Huang & Zhang, 2020; Feng et al., 2022). BASES (Cai et al., 2022b), a recent Black-box Surrogate Ensemble Search attack achieved more than 95% ASR with a few queries (<5) on average across a large scope of non-robust models.
Though the initial motivation for black-box attacks and their evaluation protocol is to improve the realism of robustness evaluations, they generally suffer from a common pitfall: they do not consider robust models which have already demonstrated their improved performance against white-box attacks. We hypothesize that the evaluation results of black-box attacks can be misleading because even simple robustification mechanisms can be sufficient to successfully evade black-box attacks. Going further, we want to assess whether innovations in defenses provide positive benefits against black-box attacks like they did in white-box settings.
@@
To the best of our knowledge, this paper is the first to study the effectiveness of black-box attacks against strandardized defenses to demonstrate the need to confront black-box attacks to existing defenses.
Our contributions can be summarized as follows:
@@
    We show that white-box robustness could be served as a proxy for black-box robustness. Defenses optimized against AutoAttack generalize well to black-box scenarios.
@@
    We demonstrate that these effective defense mechanisms can inadvertently contribute to enhancing the success rate of black-box attacks. By using robust models as surrogates, attacks can generate adversarial examples more likely to transfer to robust models, leading to an average increase in success rate of 6.42 percentage points across attacks and target models.
@@
Our contributions send a clear message to the adversarial research community regarding the importance of considering white-box defenses in the evaluation of black-box attacks. They also urge the researchers working on defense mechanisms to study how their defenses can help improve transfer-based attacks.
2 Related Work
2.1 Black-Box Attacks: Transfer-Based and Query-Based Methods
@@
Black-box attacks have seen notable advancements, especially within transfer-based and query-based approaches. Transfer-based attacks rely on adversarial examples generated by a surrogate model, which are then tested on a target model. Early approaches like Projected gradient descent (PGD) (Madry et al., 2017) laid the groundwork for such attacks by iteratively optimizing perturbations to maximize the success rate on the target model . This foundational method was enhanced by the Momentum Iterative Fast Gradient Sign Method (MI-FGSM), which incorporated momentum into PGD to improve transferability and stability across iterations (Dong et al., 2018b). Further innovations, such as the Diverse Input Fast Gradient Sign Method (DI-FGSM) and Translation-Invariant FGSM (TI-FGSM), introduced input transformations and image translations, respectively, to increase the robustness and generalizability of adversarial examples across different models (Xie et al., 2019b; Dong et al., 2019).
@@
Subsequent methods like Variance Tuning (VMI, VNI) refined gradient calculations by considering gradient variance across iterations (Wang & He, 2021), leading to more effective adversarial perturbations. Similarly, ADMIX leveraged diverse inputs by mixing the target image with randomly sampled images to generate more transferable attacks (Wang et al., 2021). Skip Gradient Method (SGM) focused on model architecture, by using skip connections which facilitate the creation of adversarial examples that have high transferability (Wu et al., 2020b). Universal Adversarial Perturbations (UAP) took a different approach by focusing on universality, generating a single small perturbation that can effectively disrupt a classifier across most natural images (Moosavi-Dezfooli et al., 2017). Notably, Ghost networks (GHOST) and Large Geometric Vicinity (LGV) contributed to transferability by modifying skip connections in surrogate models and by exploring variations in surrogate models vicinity, respectively (Li et al., 2020; Gubri et al., 2022).
@@
In query-based attacks, which generally have higher success rates than transfer attacks, methods like Zeroth Order Optimization (ZOO), Decision Boundary Attack (DBA), Natural Evolutionary Strategies (NES),and HopSkipJump rely on iterative query feedback to optimize perturbations, often not utilizing surrogate models to boost efficiency (Brendel et al., 2018; Chen et al., 2017; Ilyas et al., 2018; Chen et al., 2020). Recent work, such as BASES and Transferable Model-based Embedding (TREMBA) (Cai et al., 2022a; Huang & Zhang, 2019), integrates both surrogate representations and query feedback to create highly effective black-box attacks. BASES leverages an ensemble of surrogates with query feedback to dynamically adjust surrogate weighting, while TREMBA utilizes a generator trained with surrogate models, exploring latent space for more effective query attacks.
2.2 Defenses Against AutoAttack
@@
In parallel, adversarial defenses have evolved to mitigate the risks posed by increasingly sophisticated white-box attacks. As defenses became more robust, the need for a reliable, standardized benchmark to accurately assess their effectiveness grew increasingly apparent. In response, AutoAttack emerged as a reliable, parameter-free method for evaluating adversarial robustness, offering a computationally affordable, and a standardized benchmark applicable to various models (Croce & Hein, 2020b). Building on this, Robustbench established a leaderboard specifically for evaluating adversarial robustness, as outlined by (Croce et al., 2020). This leaderboard includes results on the ImageNet large-scale datasets (Deng et al., 2009), and offers a ranking comparison of various defense strategies across specific perturbation budgets. Among the defenses featured on the leaderboard are those proposed by (Salman et al., 2020; Singh et al., 2024; Liu et al., 2024; Bai et al., 2024; Madry et al., 2017), offering a thorough assesment on the effectiveness of the leading defense mechanisms. Traditional defenses, such as Madry’s adversarial training, initially combined clean and adversarial data during training, establishing a baseline for robustness (Madry et al., 2017). Recent advancements have incorporated complex pre-training schemes and data augmentations (e.g., ConvStem architectures, RandAugment, MixUp, and CutMix). Singh et al. (2024) made alterations to the convnets architecture, particularly substituting PatchStem with ConvStem, and modified the training scheme to enhance the robustness against previously unencountered l1 and l2 threat models. Liu et al. (2024) further advances the training scheme with large-scale pre-training, combined with label smoothing, weight decay, and Exponential Moving Averaging (EMA) to enhance generalization against unseen adversarial examples. Bai et al. (2024) builds on a giving defense framework (Liu et al., 2024), focusing specifically on balancing clean and robust accuracy through the nonlinear mixing of robust and standard models, addressing the often-seen tradeoff between clean and robust performance.
2.3 Robustness Under Black-Box Setting
@@
In this part, we focus on how our work introduces a novel contribution compared to previous studies on defenses against black-box adversarial attacks in image classification tasks. We discuss four representative works (Dong et al., 2020; Mahmood et al., 2021a; b; Ghaffari Laleh et al., 2022), highlighting their key contributions and how our approach builds upon or complements these efforts by addressing gaps and introducing new perspectives in this domain.
@@
Dong et al. (2020) evaluated adversarial robustness in image classification tasks against white-box and black-box attacks. Their work addressed various defense techniques such as robust training, input transformation, randomization, certified defenses, and model ensembling. Our study seeks to complement this by focusing specifically on SOTA robust training-based defenses. Moreover, our work explores a broader landscape of black-box attacks by including recent methods such as ADMIX, BASES, and TREMBA. In doing so, we provide an analysis of leading robust training defenses against recent black-box adversarial strategies.
@@
Mahmood et al. (2021a) noted that the majority of existing defenses primarily address white-box attacks, neglecting the crucial aspect of black-box adversarial robustness. They provided a wide evaluation of adversarial defenses with a Convolutional neural network (CNN) architecture on benchmark datasets such as CIFAR-10 (Krizhevsky, 2009) and Fashion-MNIST (Xiao et al., 2017) against 12 attacks. In a follow-up work, Mahmood et al. (2021b) extended this investigation to Vision Transformers (ViTs), evaluating their adversarial robustness against white-box attacks and two black-box attacks on ImageNet. They suggested that ensemble defenses can enhance robustness when attackers lack access to model gradients. Additionally, They examined the transferability of adversarial examples between CNNs and transformers. Building upon and complementing these efforts, we conduct our evaluations on ImageNet, incorporating both CNN and transformer architectures, and including an ensemble defense strategy that combines robust and standard base models. Our approach further investigates transferability between adversarially trained and standardly trained models. We leverage SOTA robust models to challenge the attacks used to evaluate robustness.
@@
In the field of computational pathology, Ghaffari Laleh et al. (2022) examined adversarial robustness with a focus on CNN and transformer models. This work addressed certain adversarial attacks (including Fast Adaptive Boundary (FAB) (Croce & Hein, 2020a), and Square (Andriushchenko et al., 2020)), as well as adversarial training using (PGD) with Dual-Batch Normalization (DBN). Our work extends this line of inquiry by incorporating SOTA adversarially trained models and a more comprehensive suite of black-box attacks, providing additional context for these attacks by increasing their number and offering a more challenging confrontation against SOTA defenses.
3 Success Rate of Black-Box Attacks Against Adversarial Training

### v2

@@ Abstract
Although adversarial robustness has been extensively studied in white-box settings, recent advances in black-box attacks (including transfer- and query-based approaches) are primarily benchmarked against weak defenses, leaving a significant gap in the evaluation of their effectiveness against more recent and moderate robust models (e.g., those featured in the Robustbench leaderboard). In this work, we argue that this gap is problematic and the previous benchmarks conclusions do not hold under more robust evaluation frameworks, leading to contradicting conclusions on the transferability of adversarial attacks. We extensively evaluate the effectiveness of 13 popular black-box attacks, representative of the top ten popular transferability theories. We benchmark these attacks against eight top-performing and standard defense mechanisms on the ImageNet dataset. Our empirical evaluation reveals the following key findings: (1) the most advanced black-box attacks struggle to succeed even against simple adversarially trained models; (2) robust models that are optimized to withstand strong white-box attacks, such as AutoAttack, also exhibit enhanced resilience against black-box attacks; and (3) robustness alignment between the surrogate models and the target model can significantly impact the success rate of transfer-based attacks.
Index Terms:
adversarial attack, robustbench, defenses.
[…]
I Introduction
TABLE I: Theories about the transferability of adversarial examples. In bold when the evaluation protocol of the black-box is sufficiently robust. We report the largest image dataset, the smallest perturbation distance, and the deep learning model families used in the evaluation. FNN for feed-forward neural network, CNN for convolutional neural network
Theory Dataset Representative attacks Attack budgets Target architectures Target defenses
Linearity [27] ImageNet FGSM $L_{2}=0.25$ Convolution FGSM AT
Decision Boundary Similarity [44] ImageNet FGSM $L_{0}=23$ Convolution $\varnothing$
Gradient Similarity [19] MNIST FGSM $L_{2}=1$ FNN $\varnothing$
Adversarial Subspace [68] MNIST FGSM $L_{2}=4$ FNN, CNN FGSM AT
Non-Robust Features [37] ImageNet PGD $L_{2}=3$ CNN PGD AT
Model complexity [76] ImageNet IFGSM $L_{2}=15$ CNN $\varnothing$
Perturbation interactions [71] ImageNet PGD $L_{\infty}=16/255$ CNN FGSM AT
Spectral properties [14] CIFAR-10 PGD $\mathbf{L_{\infty}=2/255}$ FNN, CNN PGD AT
Knowledge transferability [42] STL-10 PGD $\mathbf{L_{2}=2/255}$ CNN $\varnothing$
Generalization gap [72] CIFAR-100 PGD $L_{\infty}=8/255$ CNN $\varnothing$
Proposed protocol ImageNet AutoAttack, BASES $\mathbf{L_{\infty}=4/255}$ CNN, Transformers Robustbench
@@ I Introduction
Since the discovery that deep learning models are susceptible to minor input disturbances, resulting in adversarial examples (AE) [27], the development of robust models has become one of the most active topics in the machine learning community. This topic has been thoroughly explored for white-box settings, settings where attackers have full knowledge of the target, leading to powerful standardized attacks [18] and benchmarks [15].
In parallel, more ”realistic” studies of adversarial vulnerability involved restricting the knowledge or the capabilities of the attacker. These scenarios are referred to as black-box scenarios (or gray-box when partial knowledge of the target is available [30]). State-of-the-art (SoTA) black-box attacks leverage one or both of the following mechanisms: (1) adversarial examples transferability and (2) iterative queries with meta-heuristics.
@@
Transferability methods craft adversarial examples on a specifically built surrogate model (with white-box access) such that this example should transfer to (i.e. also fools) a target model (to which they have black-box access). Iterative query methods use query feedback from the target model to guide or refine the optimization process to uncover its vulnerabilities.
Black-box attacks are powerful tools to study real-world attack scenarios and to better understand the mechanisms behind adversarial vulnerabilities of deep learning models, in particular through the study of attacks’ transferability. However, they still rely on outdated and unstandardized evaluation practices (cf. Table I). In contrast to white-box research, recent black-box attacks are still designed and evaluated using large epsilon noise budgets, small architectures, and non-robust target models.
Traditional theories such as the decision boundary similarity [44], the adversarial subspace theory [68], and the non-robust features [37] are still well recognized, but have been demonstrated using simple transfer attacks, with very large budgets and non-robust models. In contrast, Robustbench [15] recommended much stronger models and enforced a budget of $\epsilon=4/255$ as a reasonable evaluation budget.
At least ten of the main theories have been empirically demonstrated using incomplete black-box attacks protocols (Table I): linearity [27], gradient similarity [19], adversarial subspace [68], non-robust features [37], model complexity [76], adversarial perturbation interactions [71], spectral properties [14], knowledge transferability [42], generalization gap [72], and decision boundary similarity [44, 34].
Our work does not intend to challenge all previous work and theories about the vulnerability of deep learning models to adversarial attacks; indeed, many theories were evaluated with the best protocols at that time. We would rather like to demonstrate that adhering to more stringent evaluation protocols leads to different insights and conclusions on the vulnerability of the models and effectiveness of the black-box attacks.
In addition, given the importance of black-box attacks in assessing real-world robustness of ML systems put in production, we also hypothesize that the previous evaluation results of black-box attacks can be misleading. Even old robustification mechanisms enable us to successfully evade supposedly SoTA black-box attacks. LGV [28], one of the strongest transfer attacks, sees its success rate drop from $87.39\%$ to $2.17\%$ under a 2020 defense by Salman et al.[59]. Our first insight suggests to reconsider previous claims of effectiveness of black-box attacks and demonstrates that simple adversarial training is sufficient against the majority of SoTA black-box attacks.
While our first insight challenges the effectiveness of SoTA attacks, we tackle the benefits of SoTA defenses in our second study. We evaluated 9 defenses from Robustbench white-box ImageNet benchmark [16] and categorized these defenses into four families. Our analysis demonstrates that within the same family of defenses, architectures and model size have limited impact on the effectiveness of defenses. Our main insight from this study is that SoTA defenses are starting to plateau and show diminishing returns against SoTA black-box attacks.
Finally, our work argues that an incomplete evaluation of the interactions between new adversarial defenses and black-box attacks can actually lead new defenses to strengthen already existing black-box attacks, opening new surface attacks that were not considered before. For example, BASES [8] can leverage recent robust models from Liu et al.[43] to increase its effectiveness against the most robust models from $1.26\%$ to $12.55\%$. In fact, we analyzed the top 20 defenses in the Robustbench ImageNet leaderboard ($L_{\infty}$, eps=4/255), and none have explored how their design choices might inadvertently aid black-box attacks. Even Blackboxbench, the most recent and complete benchmark of black-box attacks [85], did not consider the interactions between SoTA white-box defenses and SoTA black-box attacks. Our last insight highlights the new risks posed by SoTA defenses, that can be leveraged to improve existing (weak) attacks.
@@
To the best of our knowledge, this paper is the first to study the effectiveness of black-box attacks against standardized protocols and to demonstrate the need to confront black-box attacks to robust defenses and up-to-date evaluation settings. Our contributions can be summarized as follows:
@@
    We show that white-box robustness could serve as a proxy for black-box robustness. Defenses optimized against AutoAttack generalize well to black-box scenarios.
@@
    We demonstrate that these effective defense mechanisms can inadvertently contribute to enhancing the success rate of black-box attacks. By using robust models as surrogates, attacks can generate adversarial examples more likely to transfer to robust models, leading to an average increase in success rate of 6.49 percentage points across attacks and target models.
@@
Our contributions send a clear message to the adversarial research community on the importance of considering white-box defenses when designing or evaluating black-box attacks, or exploring transferability mechanisms and theories. They also urge researchers working on defense mechanisms to consider how their recent defenses can help malicious parties improve existing black-box attacks, and study how that their work could lead to stronger attacks against seemingly robust defenses. We provide a replication package with our experiments and benchmark on https://figshare.com/projects/RobustBlack/265786.
II Background
II-A Preliminaries
Adversarial perturbation [15]
Let $x\in\mathbb{R}^{d}$ be an input point and $y\in\{1,\ldots,C\}$ be its correct label. For a classifier $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$, we define a successful adversarial perturbation with respect to the perturbation set $\Delta\subseteq\mathbb{R}^{d}$ as a vector $\vec{\delta}\in\mathbb{R}^{d}$ such that
$\mathop{\rm arg\,max}\limits_{c\in\{1,\ldots,C\}}f({x}+\delta)_{c}\neq y\quad\textrm{and}\quad{\delta}\in\Delta,$
where the perturbation set $\Delta$ is chosen such that all examples in $x+\delta$ have $y$ as their true label. This motivates a robustness measure called robust accuracy, which is the fraction of data points on which the classifier $f$ predicts the correct class for all possible perturbations from the set $\Delta$. Computing the exact Robust Accuracy (RA) is in general intractable and, when considering $\ell_{p}$-balls as $\Delta$, NP-hard even for single-layer neural networks. In practice, an upper bound for robust accuracy is determined by adversarial attacks, generally involving optimization of a differentiable loss function or a reward through search algorithms that aim to identify a successful adversarial perturbation. The tightness of the upper bound depends on the strength of the attack.
Success rate
Given that different models show varying clean performances (e.g., test accuracy on the original set), robust accuracy will be impacted by the initial performance as much as the intrinsic robustness of the models. Thus, we base our study on an agnostic test performance metric: attack success rate (ASR). We define ASR for a classifier $f$ under attack $\mathcal{A};\mathcal{A}(x)={x}+\delta$ as
$\text{ASR}(f,\mathcal{A})=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathbb{I}\left(f(\mathcal{A}(x))\neq f(x)\right)\right].$
For a model with a 100% test accuracy, the two metrics are related $\text{ASR}=1-\text{RA}$.
II-B Black-Box Attacks: Transfer-Based and Query-Based Methods
@@
Black-box attacks have seen notable advances, especially within transfer-based and query-based approaches. Transfer-based attacks rely on adversarial examples generated by a surrogate model, which are then tested on a target model. Early approaches like Projected Gradient Descent (PGD) [47] laid the groundwork for such attacks by iteratively optimizing perturbations to maximize the success rate on the target model. This foundational method was enhanced by the Momentum Iterative Fast Gradient Sign Method (MI-FGSM), which incorporated momentum into PGD to improve transferability and stability across iterations [22]. Further innovations, such as the Diverse Input Fast Gradient Sign Method (DI-FGSM) and Translation-Invariant FGSM (TI-FGSM), introduced input transformations and image translations, respectively, to increase the robustness and generalizability of adversarial examples across different models [79, 23].
@@
Subsequent methods such as Variance Tuning (VMI, VNI) refined gradient calculations considering gradient variance between iterations [69], leading to more effective adversarial perturbations. Similarly, ADMIX leveraged diverse inputs by mixing the target image with randomly sampled images to generate more transferable attacks [70]. The Skip Gradient Method (SGM) focused on model architecture, using skip connections which facilitate the creation of adversarial examples that have high transferability [75]. Universal Adversarial Perturbations (UAP) took a different approach by focusing on universality, generating a single small perturbation that can effectively disrupt a classifier in most natural images [51]. In particular, Ghost networks (GHOST) and Large Geometric Vicinity (LGV) contributed to transferability by modifying skip connections in surrogate models and by exploring variations in surrogate models’ vicinity, respectively [41, 28].
@@
In query-based attacks, which generally have higher success rates than transfer attacks [2], methods like Zeroth Order Optimization (ZOO) [6], Decision Boundary Attack (DBA) [36], Natural Evolutionary Strategies (NES) [9], HopSkipJump [9], SignFlip [12], and Square [3] rely on iterative query feedback to optimize perturbations, often not utilizing surrogate models to boost efficiency .
The following attacks then focused on reducing the number of queries. [2] introduced a query-efficient black-box attack using randomized square-shaped updates at image boundaries, SimBA [29] proposed a simple attack using orthogonal search directions (e.g., DCT basis). Sign-OPT [13] introduces a hard-label attack estimating gradient signs instead of magnitudes, while SignHunt [1] uses sign bits for gradient estimation. to minimize the number of queries. RayS [10] reformulates the boundary search as discrete optimization, eliminating gradient estimation.
Recent work combined both surrogate representations and query feedback to create highly effective and efficient black-box attacks. Representative attacks include Transferable Model-based Embedding (TREMBA) and BASES [33, 7]. BASES leverages an ensemble of surrogates with query feedback to dynamically adjust surrogate weighting, while TREMBA utilizes a generator trained with surrogate models, exploring latent space for more effective query attacks.
II-C Defenses Against Strong Adversarial Attacks
@@
In parallel, adversarial defenses have evolved to mitigate the risks posed by increasingly sophisticated white-box attacks. As defenses became more robust, the need for a reliable, standardized benchmark to accurately assess their effectiveness grew increasingly apparent. In response, AutoAttack emerged as a reliable, parameter-free method to evaluate adversarial robustness, offering a computationally affordable and a standardized benchmark applicable to various models [18]. Building on this, Robustbench established a leaderboard specifically to evaluate adversarial robustness, as outlined by [15]. This leaderboard includes results on ImageNet large-scale datasets [20], and offers a ranking comparison of various defense strategies in specific perturbation budgets. Among the defenses featured on the leaderboard are those proposed by [59, 61, 43, 5, 47], which offers a thorough assessment of the effectiveness of the leading defense mechanisms. Traditional defenses, such as Madry’s adversarial training, initially combined clean and adversarial data during training, establishing a baseline for robustness [47]. Recent advancements have incorporated complex pre-training schemes and data augmentations (e.g., ConvStem architectures, RandAugment, MixUp, and CutMix). [61] made alterations to the convnets architecture, particularly substituting PatchStem with ConvStem, and modified the training scheme to enhance the robustness against previously unencountered l1 and l2 threat models. [43] further advances the training scheme with large-scale pre-training, combined with label smoothing, weight decay, and Exponential Moving Averaging (EMA) to enhance generalization against unseen adversarial examples. [5] builds on a given defense framework [43], focusing specifically on balancing clean and robust accuracy through the nonlinear mixing of robust and standard models, addressing the often-seen tradeoff between clean and robust performance.
II-D Robustness Under Black-Box Setting
@@
In this part, we focus on how our work introduces a novel contribution compared to previous studies on defenses against black-box adversarial attacks in image classification tasks. We discuss the following representative works [53, 21, 48, 49, 26, 85], highlighting their key contributions and how our approach builds on or complements these efforts by addressing gaps and introducing new perspectives in this domain.
[53] introduced a black-box attack strategy targeting deep neural network (DNN) models and evaluated its effectiveness against adversarial training and defensive distillation. They varied the magnitude of the perturbation used during training and the attack phase and demonstrated that small perturbations of adversarial training in training led to gradient masking, which their attack could bypass by doubling the perturbation budget in the attacking phase, whereas larger perturbations in training improved robustness. In contrast, our work focuses on evaluating adversarially trained SOTA defenses against black-box attacks using a fixed small perturbation 4 over 255 in the attack phase.
@@
[21] evaluated adversarial robustness in image classification tasks against white-box and black-box attacks. Their work addressed various defense techniques such as robust training, input transformation, randomization, certified defenses, and model ensembling. Our study seeks to complement this by focusing specifically on SoTA robust training-based defenses. Moreover, our work explores a broader landscape of black-box attacks by including recent methods such as ADMIX, BASES, and TREMBA. In doing so, we provide an analysis of the leading robust training defenses against recent black-box adversarial strategies.
@@
[48] noted that the majority of existing defenses primarily address white-box attacks, neglecting the crucial aspect of black-box adversarial robustness. They provided a wide evaluation of adversarial defenses with a Convolutional Neural Network (CNN) architecture on benchmark datasets such as CIFAR-10 [40] and Fashion-MNIST [77] against 12 attacks. In a follow-up work, [49] extended this investigation to Vision Transformers (ViTs), evaluating their adversarial robustness against white-box attacks and two black-box attacks on ImageNet. They suggested that ensemble defenses can enhance robustness when attackers lack access to model gradients. Additionally, they examined the transferability of adversarial examples between CNNs and transformers. Building upon and complementing these efforts, we conduct our evaluations on ImageNet, incorporating both CNN and transformer architectures, and including an ensemble defense strategy that combines robust and standard base models. Our approach further investigates transferability between adversarially trained and standardly trained models. We leverage SoTA robust models to challenge the attacks used to evaluate robustness.
@@
In the field of computational pathology, [26] examined adversarial robustness with a focus on CNN and transformer models. This work addressed certain adversarial attacks (including Fast Adaptive Boundary (FAB) [17], and Square [4]), as well as adversarial training using (PGD) with Dual-Batch Normalization (DBN). Our work extends this line of inquiry by incorporating adversarially trained SOTA models and a more comprehensive suite of black-box attacks, providing additional context for these attacks by increasing their number and offering a more challenging confrontation against SoTA defenses.
Zheng et al.[85] presented the BlackboxBench benchmark, which offers a comprehensive study for evaluating black-box adversarial attacks. Specifically, they implemented 59 black-box attack methods, including 29 query-based and 30 transfer-based algorithms. They conducted evaluations on two datasets (CIFAR-10 and ImageNet) and examined four attack settings (targeted and untargeted attacks under $L_{2}$ and $L_{\infty}$ norms). They explored both regular and adversarially trained models . Our work expands the evaluation of target and surrogate models by assessing black-box attacks with the most recent SoTA adversarially trained models from the RobustBench leaderboard. We revisit the role of alignment between surrogate and target models, confirming that training-strategy compatibility enhances transferability, a trend also observed in BlackboxBench, where we updated this analysis with SoTA defenses both as surrogates and as targets.
A recent line of research investigated preprocessing and randomness injection defenses against black-box attacks. [55] proposes adding Gaussian noise to queries to disrupt gradient estimation in black-box attacks, combining it with Gaussian-augmentation fine-tuning for improved robustness-accuracy trade-offs; [11] proposed the Adversarial Attack on Attackers (AAA) defense. It introduces output logit perturbation to misdirect score-based attacks while preserving clean accuracy and improving calibration. [52] showed that randomizing hidden features provides better robustness than input randomization against query-based attacks, with plug-and-play implementation and minimal accuracy loss. [62] demonstrated that unaware preprocessing reduces attack efficacy by 7×, and designed preprocessor-aware attacks that easily overcome such defenses.
II-E Explaining Transferability and Robustness in Black-Box Setting
Various research investigated the factors behind the transferability (or lack-off) of adversarial examples. In our study, we covered each theory with at least one representative method.
Decision Boundary Similarity [44]
Models trained on the same task develop similar decision boundaries. Adversarial perturbations are highly aligned with weight vectors across models. Attacks relying on these mechanisms include PGD and HopSkipJump.
Gradient Similarity [19]
Two main factors contribute to transferability: (1) similarity between gradient directions of source and target models, and (2) smoothness of the loss landscape. Higher gradient similarity and lower variance in loss landscapes lead to increased transferability. Representative attacks include MI-FGSM and Variance Tuning (VMI/VNI).
Shared Adversarial Subspaces [68]
Adversarial examples span contiguous subspaces of large dimensionality. When different models are trained on the same task, a fraction of their adversarial subspaces overlap. DI-FGSM, Ghost Networks, and LGV use these mechanisms.
Non-Robust Features [37]
Models learn both robust features (perceivable by humans) and non-robust features (imperceptible but statistically correlated with labels). Adversarial examples exploit non-robust features. ADMIX and NES leverage this theory.
Model Complexity/Capacity [76]
Model-specific factors, including architecture and capacity, influence transferability. Adversarial examples generated from simpler/shallower models tend to transfer better to complex models (e.g. SGM).
Interaction-Based [71]
There is a negative correlation between adversarial transferability and interactions inside adversarial perturbations. Less interaction leads to higher transferability. TI-FGSM falls into this category.
Knowledge Transferability [42]
Models with high knowledge transfer (e.g., via fine-tuning) exhibit stronger adversarial example transfer. The embeddings trained in TREMBA leverage this knowledge transfer.
Flatness/Manifold [25]
There are conflicting results that higher flatness of adversarial examples enables better cross-model transferability. BASES, through its search algorithm, finds optimal weights to explore the manifold.
Other researchers specifically explored the link between white-box robustness and transferability. [63] investigated the link between the robustness of the source models and the robustness of the target models for transferable attacks. Their results show as expected that adversarial examples generated using non-robust networks do not transfer to the adversarially trained networks. However, what they consider as robust source models is a simple Resnet50 with vanilla adversarial training, which has since been shown to only be slightly robust [15]. Our study, on the other hand, leverages state-of-the-art robustification mechanisms and explores a large set of architectures for robust surrogates and targets. Although our initial results match, our study highlights further insights (e.g., using robust sources is actually detrimental when targeting non-robust targets).
[84] confirmed the impact of model smoothness and gradient similarity by exploring the impact of weakly adversarial training (that is, training with mildly perturbed examples). The approach used to train the model, including data augmentation, synthetic data, regularization are considered in some of the models of our study. Our work focuses on robustification mechanism using extreme augmentations. All in all, our results are complementary; as they explore ”mildly robust models” and we focus on ”extremely robust models”.
III Success Rate of Black-Box Attacks Against Adversarial Training